OpenStudy (anonymous):

5 powered x-3 = 137 solve

OpenStudy (anonymous):

5^(x-3)=137?

OpenStudy (anonymous):

\[ 5^{(x-3)} = 137 \] \[ (x-3) \times log 5 = log 137 \] \[ (x-3) = \frac{log 137}{log5} \] \[ x = 3 + \frac{log 137}{log5} \]

OpenStudy (anonymous):

\[5^{x-3}=137\] \[x-3=\log_5(137)\] \[x=\log_5(137)+3\]